OpenStudy (anonymous):

graph the line containing the given pair of points and find the slope. (4,0) (-2,-5) How do I find the slope?

OpenStudy (anonymous):

m=(y1-y)/(x1-x)=(0-(-5))/(4-(-2))

OpenStudy (anonymous):

m=5/6

OpenStudy (mathteacher1729):

This might make the formula a little easier to understand. Basically you're measuring the "rise" and dividing that by the "run". (see pic)

OpenStudy (anonymous):

OOOOHHHHH well that makes it very easy lol thanks you two!!!